OpenStudy (anonymous):

Put the equation in slope-intercept form and identify the slope of the line. 3x + 9y = 12

OpenStudy (anonymous):

Perfect! Want yo try to do it by yourself? ( y =ax + b)

OpenStudy (anonymous):

D.) None of the Above

OpenStudy (anonymous):

ok hang on

OpenStudy (anonymous):

Remember, you're trying to make y by itself.

OpenStudy (anonymous):

so is it B

OpenStudy (anonymous):

If by B you mean -2/3 then no

OpenStudy (anonymous):

yes that what I meant lol. Darn...

OpenStudy (anonymous):

Already from the start you can see that they can be divided by 3.

OpenStudy (anonymous):

So it becomes x + 3y = 4 Is it easier to solve now?

OpenStudy (anonymous):

i got a decimal again when i 3/4

OpenStudy (anonymous):

Wait, don't think of this yet. Bring your x over to the other side. What is the new equation now?

OpenStudy (anonymous):

3y=4x?

OpenStudy (anonymous):

You substract the x. Don't multiply.

OpenStudy (anonymous):

You only multiply when something is divided. (ex: 1/x = 2 then it becomes 2x=1)

OpenStudy (anonymous):

so its just 3y=4?

OpenStudy (anonymous):

x-x+3y = 4-x 3y = 4-x

OpenStudy (anonymous):

now what do you divide the whole equation by to make y by itself?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

Exactly!

OpenStudy (anonymous):

Can you do it?

OpenStudy (anonymous):

im kinda ify

OpenStudy (anonymous):

ify?

OpenStudy (anonymous):

not sure, thats what i meant by ify lol. sorry :P

OpenStudy (anonymous):

3y = 4-x 4/3 = -x/3 =

OpenStudy (anonymous):

y = -1/3x +4/3

OpenStudy (anonymous):

The slope is -1/3

OpenStudy (anonymous):

i appreciate it!!